OpenStudy (anonymous):

Find the midpoint of segment P Q if P(10, 2) and Q(8, 0)? Answer (-9, -1) (-1, 9) (1, 9) (9, 1)

hartnn (hartnn):

same formula! The co-ordinates of midpoint(x,y) on the line segment with end-points (x1,y1) and (x2,y2) is just the average of co-ordinates, given by : \(\large x=\frac{x_1+x_2}{2}\) \(\large y=\frac{y_1+y_2}{2}\) now,just put the values and find the co-ordinates of midpoint (x,y).

hartnn (hartnn):

what is (10+8)/2 =... ?

OpenStudy (anonymous):

9

hartnn (hartnn):

yes! what is (2+0)/2 =...

OpenStudy (anonymous):

1

hartnn (hartnn):

so, its (9,1) :)